OpenStudy (anonymous):

can some1 check my answer

OpenStudy (anonymous):

Find the volume of the prism. 5m 4m 3m A. 36 m3 B. 60 m3 C. 72 m3 D. 120 m3 i pick B

OpenStudy (anonymous):

@YoungStudier help

jhonyy9 (jhonyy9):

so where is your answer ?

OpenStudy (anonymous):

i pick B

jhonyy9 (jhonyy9):

you are right but can you tell me how you calculi volume of a rectangular prism ?

OpenStudy (anonymous):

i did 3*4*5 = 60

jhonyy9 (jhonyy9):

yes area of base time height nice

OpenStudy (anonymous):

Find the volume of the prism. https://static.k12.com/calms_media/media/559500_560000/559718/2/d278604ccddd1d1134455af80482979bcf91bc61/67447_question.jpg A. 20 yd3 B. 50 yd3 C. 200 yd3 D. 250 yd3

OpenStudy (anonymous):

i pick C.

jhonyy9 (jhonyy9):

so rewrite it please in question cuolomn - ok .?

jhonyy9 (jhonyy9):

but C. is right sure

OpenStudy (anonymous):

A box containing tissues has the shape of a rectangular prism measuring 19 inches by 23 inches by 8 inches. What is the volume of the box? A. 3496 in3 B. 1748 in3 C. 300 in3 D. 50 in3 ~i pick A. for this one

jhonyy9 (jhonyy9):

so than you have got it from 19*23*8 so than this is right sure

OpenStudy (anonymous):

thx

jhonyy9 (jhonyy9):

was my pleasure

OpenStudy (anonymous):

A candy bar is packaged in a box that has the shape of a triangular prism. The triangular ends of the box each has a base 4 cm long and a height that is 3 cm long. The box is 15 cm long. What is the volume of the box? A. 22 cm3 B. 44 cm3 C. 90 cm3 D. 180 cm3 ~~~~~~~~~~~i pick D~~~~~~~~~~~~~~~

jhonyy9 (jhonyy9):

do you understand the text of this exercise sure right ?

jhonyy9 (jhonyy9):

there is a triangular prism - yes ? so than how you calculi area of a triangle ?

jhonyy9 (jhonyy9):

so you dont know formula for area of a triangle ?

OpenStudy (anonymous):

no :/

jhonyy9 (jhonyy9):

but you have learned it sure - yes ?

OpenStudy (anonymous):

yea

jhonyy9 (jhonyy9):

base time height divide two = area of a triangle do you remember it now ?

OpenStudy (anonymous):

yea

jhonyy9 (jhonyy9):

can using this to calculi area of base of this box ?

OpenStudy (anonymous):

i can try

jhonyy9 (jhonyy9):

ok. wait

jhonyy9 (jhonyy9):

what you know there ? that the base is 4 cm long and height is 3 cm - so than area = ?

jhonyy9 (jhonyy9):

above is formula

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

so than ? wait

jhonyy9 (jhonyy9):

area will be = ?

OpenStudy (anonymous):

180?

jhonyy9 (jhonyy9):

A=4*3/2 = ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

come please - volum = ?

jhonyy9 (jhonyy9):

A= 4*3/2 = 6 the box is long 15 cm so v= 6*15 = ?

OpenStudy (anonymous):

ok so it is 90

jhonyy9 (jhonyy9):

yes is right - do you understand it now how you solve the same exercises ?

OpenStudy (anonymous):

yea thx

jhonyy9 (jhonyy9):

ok. was my pleasure

OpenStudy (anonymous):

What is the volume of the triangular prism? https://static.k12.com/calms_media/media/1394500_1395000/1394536/1/b72009c76dcc09a2762824ae1e6f1a5a409053a1/67450_FGA_U5_L3_V3_P.jpg A. 1500 ft3 B. 2500 ft3 C. 3000 ft3 D. 4500 ft3 i pick C

OpenStudy (anonymous):

gtg bye thx

jhonyy9 (jhonyy9):

sorry i was eating a little - are you here ?

jhonyy9 (jhonyy9):

wait you to help - ok ?